OpenStudy (anonymous):

J= ∫x^2yds. it's a straight line with coordinates A(0,1) and B(3,2)

OpenStudy (anonymous):

\[\int\limits_{[AB]} x^2y ds\]

OpenStudy (anonymous):



OpenStudy (amistre64):

ds refers to a distance integral:\[\sqrt{(x')^2+(y')^2}{}\]

OpenStudy (amistre64):

might want to define your line as an equation tho

OpenStudy (anonymous):

i used a formula for the equation of the line but.. i don't think that the formula that i've got is ok. i have to make it into a double integral?

OpenStudy (amistre64):

the interval a to b defines "t" x = 0+3t ; x^2 = 9t^2 y = 1 + t

OpenStudy (amistre64):

x' = 3 y' = 1

OpenStudy (amistre64):

\[\int_{0}^{1}9t^2(1+t)~\sqrt{3^2+1^2}~dt\]

OpenStudy (anonymous):

thank you :)

OpenStudy (amistre64):

yep